OpenStudy (anonymous):

f(x)=9log(x + 10) - 9 is

OpenStudy (anonymous):

f(x)= 3log(x + 5) - 2 0 = 3log(x + 5) - 2 2 = 3log(x + 5) 2/3 = log(x + 5) (x + 5) = 10^(2/3) x = 10^(2/3) - 5 this is an exact value.